# Pre-training Cross-lingual Open Domain Question Answering with Large-scale Synthetic Supervision

### v1

@@ 5 Related Work
Multilingual Dense Retrieval
@@
Multilingual Retrieval Pre-training
Large-scale unsupervised retrieval pre-training has been shown to significantly enhance dense retrievers (Gao and Callan, 2021; Izacard et al., 2022) in processing English texts. Pre-training has also been explored in cross-lingual and multilingual dense retrieval, with a particular emphasis on augmenting the cross-lingual alignment capabilities of models. LAPCA (Abulkhanov et al., 2023) is trained through extensive cross-lingual contrastive learning, employing texts from parallel Wikipedia pages and parallel texts generated by machine translation systems. DrDecr (Li et al., 2022) learns from English models but operates on a smaller scale and relies on supervised parallel queries. In this work, we delve into the potential of large-scale unsupervised pre-training for cross-lingual dense retrieval and show that the resulting model exhibits high efficacy, outperforming many supervised ones.
Pre-training for Retrieval-Augmented Multilingual QA
@@ 6 Conclusion
Both stages in our pre-training paradigm depend on the availability of parallel Wikipedia pages. This can pose a challenge when dealing with languages that have limited resources even in terms of monolingual texts. Our approach may fail when no language links exist between English and a specific low-resource language. One may resort to employing a multi-hop approach to discover parallel Wikipedia pages, by first searching for the language linked to the low-resource language within Wikipedia and then repeating this process iteratively until reaching the corresponding English page. Another option could be relying on the generalisation of the multilingual model by training it in closely-related languages. Our analysis has revealed that incorporating a high-resource language in the pre-training phase consistently results in improvements for other languages within the same language family, which makes this issue less of a concern. Nevertheless, it remains imperative to explore methods for reducing the reliance on parallel Wikipedia texts, as this is essential to scale our method to more diverse and unique languages, which is worth exploring as a future work.
@@
  • Izacard and Grave (2021) Gautier Izacard and Edouard Grave. 2021. Leveraging passage retrieval with generative models for open domain question answering. In Proceedings of the 16th Conference of the European Chapter of the Association for Computational Linguistics: Main Volume, pages 874–880, Online. Association for Computational Linguistics.
@@
  • Zhang et al. (2021) Xinyu Zhang, Xueguang Ma, Peng Shi, and Jimmy Lin. 2021. Mr. TyDi: A multi-lingual benchmark for dense retrieval. In Proceedings of the 1st Workshop on Multilingual Representation Learning, pages 127–137, Punta Cana, Dominican Republic. Association for Computational Linguistics.
@@ A.1.1 Parallel Queries Mining
Рис. 5: The number of mined parallel queries for each language pair en-x.
@@ A.1.2 Query Transformation
We use ChatGPT to generate 32 meta-examples. We then employ LLaMA-2-7B [^9] https://huggingface.co/meta-llama/Llama-2-7b for query transformation by randomly sampling 3 meta-examples to construct prompts for each test instance, with the format as shown in Prompts D, D, D, D, D, D, and D. We use Bloomz-7B [^10] https://huggingface.co/bigscience/bloomz-7b1 for Telugu as we find LLaMA-2-7B does not work well in this language. The Question word wh_word is chosen based on the entity type of the answer according to the heuristic rules in Table 7. Ultimately, 146K examples are generated per language, resulting in a total of 1M training instances.
@@ A.2 Datasets
  • XOR-Retrieve.  This dataset is under the MIT License for non-commercial research purposes. It contains 15250 QA pairs for training and takes the 20190201 English Wikipedia dump which contains 18M passages as the retrieval database.
  • XOR-Full.  It is under the MIT License for non-commercial research purposes. It contains 61360 QA pairs for training and a collection of 43M passages as the retrieval database, which is collected from 20190201 Wikipedia dumps across 13 languages, namely English, Arabic, Finnish, Japanese, Korean, Russian, Bengali, Telugu, Indonesian, Thai, Hebrew, Swedish, and Spanish.
  • Natural Questions.  It is under the Apache License and contains 79168 QA pairs.
  • MKQA.  It is under the Apache License. This dataset covers 26 linguistically diverse languages, namely Arabic, Danish, German, English, Spanish, Finnish, French, Hebrew, Hungarian, Italian, Japanese, Korean, Khmer, Malay, Dutch, Norwegian, Polish, Portuguese, Russian, Swedish, Thai, Turkish, Vietnamese, Chinese (Simplified), Chinese (Hong Kong), and Chinese (Traditional). For the cross-lingual retrieval task, each language contains 6620 questions and the retrieval database consists of 18M English Wikipedia passages. For the multilingual QA task, each language contains 6758 questions and it uses the same retrieval database as XOR-Full.
  • Mr.TyDi.  It is under the Apache License. This dataset contains 11 languages, namely Arabic, Bengali, English, Finnish, Indonesian, Japanese, Korean, Russian, Swahili, Telugu, and Thai. The queries for testing are 1081, 111, 744, 1254, 829, 720, 421, 995, 670, 646, and 1190 in each language, respectively. The corpus sizes for retrieval are 2M, 300K, 33M, 1.9M, 1.5M, 7M, 1.5M, 9.6M, 137K, 548K, and 569K in each language, respectively.

### v3

@@ 5 Related Work
Multilingual Dense Retrieval.
@@
Multilingual Retrieval Pre-training.
Large-scale unsupervised retrieval pre-training has significantly enhanced dense retrievers (Gao and Callan, 2021; Izacard et al., 2022) in processing English texts. Pre-training has also been explored in cross-lingual and multilingual dense retrieval, with a particular emphasis on augmenting the cross-lingual alignment capabilities of models. LAPCA (Abulkhanov et al., 2023) is trained through extensive cross-lingual contrastive learning, employing texts from parallel Wikipedia pages and parallel texts generated by machine translation systems. DrDecr (Li et al., 2022) learns from English models but operates on a smaller scale and relies on supervised parallel queries. In this work, we delve into the potential of large-scale unsupervised pre-training for cross-lingual dense retrieval and show that the resulting model exhibits high efficacy, outperforming many supervised ones.
Pre-training for Retrieval-Augmented Multilingual QA.
@@ 6 Conclusion
Both stages in our pre-training paradigm depend on the availability of parallel Wikipedia pages. This can pose a challenge when dealing with languages that have limited resources even in terms of monolingual texts. Our approach may fail when no language links exist between English and a specific low-resource language. One may resort to employing a multi-hop approach to discover parallel Wikipedia pages, by first searching for the language linked to the low-resource language within Wikipedia and then repeating this process iteratively until reaching the corresponding English page. Another option could be relying on the generalisation of the multilingual model by training it in closely-related languages. Our analysis has revealed that incorporating a high-resource language in the pre-training phase consistently results in improvements for other languages within the same language family (Figure 5), which makes this issue less of a concern. Nevertheless, it remains imperative to explore methods for reducing the reliance on parallel Wikipedia texts, as this is essential to scale our method to more diverse and unique languages, which is worth exploring as a future work.
@@
Acknowledgement
We thank the anonymous reviewers for their helpful feedback and suggestions. The first author is supported by the Graduate Research Scholarships funded by the University of Melbourne. This work was funded by the Australian Research Council, Discovery grant DP230102775. This research was undertaken using the LIEF HPCGPGPU Facility hosted at the University of Melbourne, which was established with the assistance of LIEF Grant LE170100200.
@@ Acknowledgement
  • Dror et al. (2018) Rotem Dror, Gili Baumer, Segev Shlomov, and Roi Reichart. 2018. The hitchhiker’s guide to testing statistical significance in natural language processing. In Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 1383–1392, Melbourne, Australia. Association for Computational Linguistics.
  • Dubey et al. (2024) Abhimanyu Dubey, Abhinav Jauhri, Abhinav Pandey, Abhishek Kadian, Ahmad Al-Dahle, Aiesha Letman, Akhil Mathur, Alan Schelten, Amy Yang, Angela Fan, Anirudh Goyal, Anthony Hartshorn, Aobo Yang, Archi Mitra, Archie Sravankumar, Artem Korenev, Arthur Hinsvark, Arun Rao, Aston Zhang, Aurelien Rodriguez, Austen Gregerson, Ava Spataru, Baptiste Roziere, Bethany Biron, Binh Tang, Bobbie Chern, Charlotte Caucheteux, Chaya Nayak, Chloe Bi, Chris Marra, Chris McConnell, Christian Keller, Christophe Touret, Chunyang Wu, Corinne Wong, Cristian Canton Ferrer, Cyrus Nikolaidis, Damien Allonsius, Daniel Song, Danielle Pintz, Danny Livshits, David Esiobu, Dhruv Choudhary, Dhruv Mahajan, Diego Garcia-Olano, Diego Perino, Dieuwke Hupkes, Egor Lakomkin, Ehab AlBadawy, Elina Lobanova, Emily Dinan, Eric Michael Smith, Filip Radenovic, Frank Zhang, Gabriel Synnaeve, Gabrielle Lee, Georgia Lewis Anderson, Graeme Nail, Gregoire Mialon, Guan Pang, Guillem Cucurell, Hailey Nguyen, Hannah Korevaar, Hu Xu, Hugo Touvron, Iliyan Zarov, Imanol Arrieta Ibarra, Isabel Kloumann, Ishan Misra, Ivan Evtimov, Jade Copet, Jaewon Lee, Jan Geffert, Jana Vranes, Jason Park, Jay Mahadeokar, Jeet Shah, Jelmer van der Linde, Jennifer Billock, Jenny Hong, Jenya Lee, Jeremy Fu, Jianfeng Chi, Jianyu Huang, Jiawen Liu, Jie Wang, Jiecao Yu, Joanna Bitton, Joe Spisak, Jongsoo Park, Joseph Rocca, Joshua Johnstun, Joshua Saxe, Junteng Jia, Kalyan Vasuden Alwala, Kartikeya Upasani, Kate Plawiak, Ke Li, Kenneth Heafield, Kevin Stone, Khalid El-Arini, Krithika Iyer, Kshitiz Malik, Kuenley Chiu, Kunal Bhalla, Lauren Rantala-Yeary, Laurens van der Maaten, Lawrence Chen, Liang Tan, Liz Jenkins, Louis Martin, Lovish Madaan, Lubo Malo, Lukas Blecher, Lukas Landzaat, Luke de Oliveira, Madeline Muzzi, Mahesh Pasupuleti, Mannat Singh, Manohar Paluri, Marcin Kardas, Mathew Oldham, Mathieu Rita, Maya Pavlova, Melanie Kambadur, Mike Lewis, Min Si, Mitesh Kumar Singh, Mona Hassan, Naman Goyal, Narjes Torabi, Nikolay Bashlykov, Nikolay Bogoychev, Niladri Chatterji, Olivier Duchenne, Onur Çelebi, Patrick Alrassy, Pengchuan Zhang, Pengwei Li, Petar Vasic, Peter Weng, Prajjwal Bhargava, Pratik Dubal, Praveen Krishnan, Punit Singh Koura, Puxin Xu, Qing He, Qingxiao Dong, Ragavan Srinivasan, Raj Ganapathy, Ramon Calderer, Ricardo Silveira Cabral, Robert Stojnic, Roberta Raileanu, Rohit Girdhar, Rohit Patel, Romain Sauvestre, Ronnie Polidoro, Roshan Sumbaly, Ross Taylor, Ruan Silva, Rui Hou, Rui Wang, Saghar Hosseini, Sahana Chennabasappa, Sanjay Singh, Sean Bell, Seohyun Sonia Kim, Sergey Edunov, Shaoliang Nie, Sharan Narang, Sharath Raparthy, Sheng Shen, Shengye Wan, Shruti Bhosale, Shun Zhang, Simon Vandenhende, Soumya Batra, Spencer Whitman, Sten Sootla, Stephane Collot, Suchin Gururangan, Sydney Borodinsky, Tamar Herman, Tara Fowler, Tarek Sheasha, Thomas Georgiou, Thomas Scialom, Tobias Speckbacher, Todor Mihaylov, Tong Xiao, Ujjwal Karn, Vedanuj Goswami, Vibhor Gupta, Vignesh Ramanathan, Viktor Kerkez, Vincent Gonguet, Virginie Do, Vish Vogeti, Vladan Petrovic, Weiwei Chu, Wenhan Xiong, Wenyin Fu, Whitney Meers, Xavier Martinet, Xiaodong Wang, Xiaoqing Ellen Tan, Xinfeng Xie, Xuchao Jia, Xuewei Wang, Yaelle Goldschlag, Yashesh Gaur, Yasmine Babaei, Yi Wen, Yiwen Song, Yuchen Zhang, Yue Li, Yuning Mao, Zacharie Delpierre Coudert, Zheng Yan, Zhengxing Chen, Zoe Papakipos, Aaditya Singh, Aaron Grattafiori, Abha Jain, Adam Kelsey, Adam Shajnfeld, Adithya Gangidi, Adolfo Victoria, Ahuva Goldstand, Ajay Menon, Ajay Sharma, Alex Boesenberg, Alex Vaughan, Alexei Baevski, Allie Feinstein, Amanda Kallet, Amit Sangani, Anam Yunus, Andrei Lupu, Andres Alvarado, Andrew Caples, Andrew Gu, Andrew Ho, Andrew Poulton, Andrew Ryan, Ankit Ramchandani, Annie Franco, Aparajita Saraf, Arkabandhu Chowdhury, Ashley Gabriel, Ashwin Bharambe, Assaf Eisenman, Azadeh Yazdan, Beau James, Ben Maurer, Benjamin Leonhardi, Bernie Huang, Beth Loyd, Beto De Paola, Bhargavi Paranjape, Bing Liu, Bo Wu, Boyu Ni, Braden Hancock, Bram Wasti, Brandon Spence, Brani Stojkovic, Brian Gamido, Britt Montalvo, Carl Parker, Carly Burton, Catalina Mejia, Changhan Wang, Changkyu Kim, Chao Zhou, Chester Hu, Ching-Hsiang Chu, Chris Cai, Chris Tindal, Christoph Feichtenhofer, Damon Civin, Dana Beaty, Daniel Kreymer, Daniel Li, Danny Wyatt, David Adkins, David Xu, Davide Testuggine, Delia David, Devi Parikh, Diana Liskovich, Didem Foss, Dingkang Wang, Duc Le, Dustin Holland, Edward Dowling, Eissa Jamil, Elaine Montgomery, Eleonora Presani, Emily Hahn, Emily Wood, Erik Brinkman, Esteban Arcaute, Evan Dunbar, Evan Smothers, Fei Sun, Felix Kreuk, Feng Tian, Firat Ozgenel, Francesco Caggioni, Francisco Guzmán, Frank Kanayet, Frank Seide, Gabriela Medina Florez, Gabriella Schwarz, Gada Badeer, Georgia Swee, Gil Halpern, Govind Thattai, Grant Herman, Grigory Sizov, Guangyi, Zhang, Guna Lakshminarayanan, Hamid Shojanazeri, Han Zou, Hannah Wang, Hanwen Zha, Haroun Habeeb, Harrison Rudolph, Helen Suk, Henry Aspegren, Hunter Goldman, Ibrahim Damlaj, Igor Molybog, Igor Tufanov, Irina-Elena Veliche, Itai Gat, Jake Weissman, James Geboski, James Kohli, Japhet Asher, Jean-Baptiste Gaya, Jeff Marcus, Jeff Tang, Jennifer Chan, Jenny Zhen, Jeremy Reizenstein, Jeremy Teboul, Jessica Zhong, Jian Jin, Jingyi Yang, Joe Cummings, Jon Carvill, Jon Shepard, Jonathan McPhie, Jonathan Torres, Josh Ginsburg, Junjie Wang, Kai Wu, Kam Hou U, Karan Saxena, Karthik Prasad, Kartikay Khandelwal, Katayoun Zand, Kathy Matosich, Kaushik Veeraraghavan, Kelly Michelena, Keqian Li, Kun Huang, Kunal Chawla, Kushal Lakhotia, Kyle Huang, Lailin Chen, Lakshya Garg, Lavender A, Leandro Silva, Lee Bell, Lei Zhang, Liangpeng Guo, Licheng Yu, Liron Moshkovich, Luca Wehrstedt, Madian Khabsa, Manav Avalani, Manish Bhatt, Maria Tsimpoukelli, Martynas Mankus, Matan Hasson, Matthew Lennie, Matthias Reso, Maxim Groshev, Maxim Naumov, Maya Lathi, Meghan Keneally, Michael L. Seltzer, Michal Valko, Michelle Restrepo, Mihir Patel, Mik Vyatskov, Mikayel Samvelyan, Mike Clark, Mike Macey, Mike Wang, Miquel Jubert Hermoso, Mo Metanat, Mohammad Rastegari, Munish Bansal, Nandhini Santhanam, Natascha Parks, Natasha White, Navyata Bawa, Nayan Singhal, Nick Egebo, Nicolas Usunier, Nikolay Pavlovich Laptev, Ning Dong, Ning Zhang, Norman Cheng, Oleg Chernoguz, Olivia Hart, Omkar Salpekar, Ozlem Kalinli, Parkin Kent, Parth Parekh, Paul Saab, Pavan Balaji, Pedro Rittner, Philip Bontrager, Pierre Roux, Piotr Dollar, Polina Zvyagina, Prashant Ratanchandani, Pritish Yuvraj, Qian Liang, Rachad Alao, Rachel Rodriguez, Rafi Ayub, Raghotham Murthy, Raghu Nayani, Rahul Mitra, Raymond Li, Rebekkah Hogan, Robin Battey, Rocky Wang, Rohan Maheswari, Russ Howes, Ruty Rinott, Sai Jayesh Bondu, Samyak Datta, Sara Chugh, Sara Hunt, Sargun Dhillon, Sasha Sidorov, Satadru Pan, Saurabh Verma, Seiji Yamamoto, Sharadh Ramaswamy, Shaun Lindsay, Shaun Lindsay, Sheng Feng, Shenghao Lin, Shengxin Cindy Zha, Shiva Shankar, Shuqiang Zhang, Shuqiang Zhang, Sinong Wang, Sneha Agarwal, Soji Sajuyigbe, Soumith Chintala, Stephanie Max, Stephen Chen, Steve Kehoe, Steve Satterfield, Sudarshan Govindaprasad, Sumit Gupta, Sungmin Cho, Sunny Virk, Suraj Subramanian, Sy Choudhury, Sydney Goldman, Tal Remez, Tamar Glaser, Tamara Best, Thilo Kohler, Thomas Robinson, Tianhe Li, Tianjun Zhang, Tim Matthews, Timothy Chou, Tzook Shaked, Varun Vontimitta, Victoria Ajayi, Victoria Montanez, Vijai Mohan, Vinay Satish Kumar, Vishal Mangla, Vítor Albiero, Vlad Ionescu, Vlad Poenaru, Vlad Tiberiu Mihailescu, Vladimir Ivanov, Wei Li, Wenchen Wang, Wenwen Jiang, Wes Bouaziz, Will Constable, Xiaocheng Tang, Xiaofang Wang, Xiaojian Wu, Xiaolan Wang, Xide Xia, Xilun Wu, Xinbo Gao, Yanjun Chen, Ye Hu, Ye Jia, Ye Qi, Yenda Li, Yilin Zhang, Ying Zhang, Yossi Adi, Youngjin Nam, Yu, Wang, Yuchen Hao, Yundi Qian, Yuzi He, Zach Rait, Zachary DeVito, Zef Rosnbrick, Zhaoduo Wen, Zhenyu Yang, and Zhiwei Zhao. 2024. The llama 3 herd of models.
@@
  • Izacard and Grave (2021a) Gautier Izacard and Edouard Grave. 2021a. Distilling knowledge from reader to retriever for question answering. In International Conference on Learning Representations.
  • Izacard and Grave (2021b) Gautier Izacard and Edouard Grave. 2021b. Leveraging passage retrieval with generative models for open domain question answering. In Proceedings of the 16th Conference of the European Chapter of the Association for Computational Linguistics: Main Volume, pages 874–880, Online. Association for Computational Linguistics.
@@
  • Sorokin et al. (2022) Nikita Sorokin, Dmitry Abulkhanov, Irina Piontkovskaya, and Valentin Malykh. 2022. Ask me anything in your native language. In Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, pages 395–406, Seattle, United States. Association for Computational Linguistics.
  • Team et al. (2024) Gemma Team, Thomas Mesnard, Cassidy Hardin, Robert Dadashi, Surya Bhupatiraju, Shreya Pathak, Laurent Sifre, Morgane Rivière, Mihir Sanjay Kale, Juliette Love, Pouya Tafti, Léonard Hussenot, Pier Giuseppe Sessa, Aakanksha Chowdhery, Adam Roberts, Aditya Barua, Alex Botev, Alex Castro-Ros, Ambrose Slone, Amélie Héliou, Andrea Tacchetti, Anna Bulanova, Antonia Paterson, Beth Tsai, Bobak Shahriari, Charline Le Lan, Christopher A. Choquette-Choo, Clément Crepy, Daniel Cer, Daphne Ippolito, David Reid, Elena Buchatskaya, Eric Ni, Eric Noland, Geng Yan, George Tucker, George-Christian Muraru, Grigory Rozhdestvenskiy, Henryk Michalewski, Ian Tenney, Ivan Grishchenko, Jacob Austin, James Keeling, Jane Labanowski, Jean-Baptiste Lespiau, Jeff Stanway, Jenny Brennan, Jeremy Chen, Johan Ferret, Justin Chiu, Justin Mao-Jones, Katherine Lee, Kathy Yu, Katie Millican, Lars Lowe Sjoesund, Lisa Lee, Lucas Dixon, Machel Reid, Maciej Mikuła, Mateo Wirth, Michael Sharman, Nikolai Chinaev, Nithum Thain, Olivier Bachem, Oscar Chang, Oscar Wahltinez, Paige Bailey, Paul Michel, Petko Yotov, Rahma Chaabouni, Ramona Comanescu, Reena Jana, Rohan Anil, Ross McIlroy, Ruibo Liu, Ryan Mullins, Samuel L Smith, Sebastian Borgeaud, Sertan Girgin, Sholto Douglas, Shree Pandya, Siamak Shakeri, Soham De, Ted Klimenko, Tom Hennigan, Vlad Feinberg, Wojciech Stokowiec, Yu hui Chen, Zafarali Ahmed, Zhitao Gong, Tris Warkentin, Ludovic Peran, Minh Giang, Clément Farabet, Oriol Vinyals, Jeff Dean, Koray Kavukcuoglu, Demis Hassabis, Zoubin Ghahramani, Douglas Eck, Joelle Barral, Fernando Pereira, Eli Collins, Armand Joulin, Noah Fiedel, Evan Senter, Alek Andreev, and Kathleen Kenealy. 2024. Gemma: Open models based on gemini research and technology.
  • Thakur et al. (2023) Nandan Thakur, Jianmo Ni, Gustavo Hernández Ábrego, John Wieting, Jimmy Lin, and Daniel Cer. 2023. Leveraging llms for synthesizing training data across many languages in multilingual dense retrieval.
@@
Overview of Appendix
Our supplementary includes the following sections:
  • Section A: Experimental Settings, including implementation details, datasets, and compared baselines.
  • Section B: Full zero-shot evaluation results on MKQA.
  • Section C: Error analysis on multilingual open-domain question answering with quantitative and qualitative results.
  • Section D: Additional numeric analysis.
  • Section E: Prompts and examples for query transformation in each target language.
@@ A.1.1 Parallel Queries Mining
Рис. 6: The number of mined parallel queries for each language pair en-x.
@@ A.1.2 Query Transformation
We use ChatGPT to generate 32 meta-examples. We then employ LLaMA-2-7B [^8] https://huggingface.co/meta-llama/Llama-2-7b for query transformation by randomly sampling 3 meta-examples to construct prompts for each test instance, with the format as shown in Prompts E, E, E, E, E, E, and E. We use Bloomz-7B [^9] https://huggingface.co/bigscience/bloomz-7b1 for Telugu as we find LLaMA-2-7B does not work well in this language. The Question word wh_word is chosen based on the entity type of the answer according to the heuristic rules in Table 11. Ultimately, 146K examples are generated per language, resulting in a total of 1M training instances.
@@ A.1.3 Training Details
For the Dense Retrieval variant, we follow the same training and hyperparameter settings. The only difference is that this configuration is significantly more efficient, with training time reduced by half for multilingual QA pre-training and fine-tuning.
@@ A.2 Datasets
  • XOR-Retrieve  (Asai et al., 2021a). It is under the MIT License. It contains 15250 QA pairs for training and takes the 20190201 English Wikipedia dump which contains 18M passages as the retrieval database.
  • XOR-Full  (Asai et al., 2021a). It is under the MIT License, containing 61360 training examples and a set of 43M passages as the retrieval corpus, collected from 20190201 Wikipedia dumps across 13 languages, namely English, Arabic, Finnish, Japanese, Korean, Russian, Bengali, Telugu, Indonesian, Thai, Hebrew, Swedish, and Spanish.
  • Natural Questions  (Kwiatkowski et al., 2019). It is under the Apache License and contains 79168 QA pairs.
  • MKQA  (Longpre et al., 2021). It is under the Apache License. This dataset covers 26 linguistically diverse languages, namely Arabic, Danish, German, English, Spanish, Finnish, French, Hebrew, Hungarian, Italian, Japanese, Korean, Khmer, Malay, Dutch, Norwegian, Polish, Portuguese, Russian, Swedish, Thai, Turkish, Vietnamese, Chinese (Simplified), Chinese (Hong Kong), and Chinese (Traditional). For the cross-lingual retrieval task, each language contains 6620 questions and the retrieval database consists of 18M English Wikipedia passages. For the multilingual QA task, each language contains 6758 questions and it uses the same retrieval database as XOR-Full.